OpenStudy (anonymous):

What number is a solution of 4 - 2(x + 4) ≤ 6 but not a solution of 4 - 2(x + 4) < 6

OpenStudy (anonymous):

x=3

OpenStudy (mathmate):

The two inequalities are similar. If it is a solution of the first, and not the second, it means that equality holds. 4 - 2(x + 4)=6 -2(x+4) = 6-4=2 (x+4)=-1 x=-1-4=-5